OpenStudy (sillybutt100):

Help with simplifying radical expressions. The two questions will be in the reply section.

OpenStudy (sillybutt100):

\[\sqrt{50}-\sqrt{72}+\sqrt{98}\]and \[\sqrt{\frac{ 5 }{ 16 }}\]

OpenStudy (sillybutt100):

@StudyGurl14

OpenStudy (studygurl14):

The first one \[\sqrt{50}-\sqrt{72}+\sqrt{98}\]

OpenStudy (studygurl14):

\[\sqrt{25\times2}-\sqrt{36\times2}+\sqrt{49\times2}\] \[5\sqrt{2}-6\sqrt{2}+7\sqrt{2} = 6\sqrt{2}\]

OpenStudy (studygurl14):

the second one \[\sqrt{\frac{5}{16}}\]

OpenStudy (studygurl14):

\[\frac{\sqrt{5}}{\sqrt{16}}\] \[\frac{\sqrt{5}}{4}\]

OpenStudy (sillybutt100):

thank you i get how to solve them now